Calculate the mass in grams of 1.88 × 10^20 molecules of diarsenic pentaoxide. Enter your answer in scientific notation.
Answer #1

Answer: We can use dimensional analysis to convert 1.88 * 1020 molecules diarsenic pentaoxide to moles of diarsenic pentaoxide, to grams of diarsenic pentaoxide.

Molar mass of diarsenic pentaoxide(As2O5): 229.8 g/mol

1 mol of diarsenic pentaoxide: 6.022 * 1023 molecules

Therefore,

Molar mass of diarsenic pentaoxide = 1 mol of diarsenic pentaoxide

229.8 g/mol = 6.022 * 1023

Mass of 1.88 * 1020 molecules= molar mass * (1.88 * 1020)/1 mol of diarsenic pentaoxide

Mass of 1.88 * 1020 molecules =  229.8*(1.88 * 1020/6.022 * 1023)

Mass of 1.88 * 1020 molecules = 0.0716 gm

So, mass in grams of 1.88 * 1020 molecules of diarsenic pentaoxide is 0.0716 gm.